OpenStudy (anonymous):

What are the x– and y–intercepts of y = 2x2 – 10x + 8

OpenStudy (anonymous):

make each of them zero for eachother and then solve it to get the answer, you will get two x-intercepts for y=0 and one for y i.e when x=0

OpenStudy (anonymous):

= 2(x2 - 5x + 4) = 0 (x-4)(x-1) = 0 x interceots are 1 and 4 y intercept occurs when x = 0 = 8